A)In a study of the decomposition of ammonia on a platinum surface at 856 °C
NH3>1/2 N2 + 3/2 H2
the following data were obtained:

[NH3], M 1.12×10-2 5.60×10-3 2.80×10-3 1.40×10-3
seconds 0 1000 1.51×103 1.76×103

Hint: It is not necessary to graph these data.
(1)The observed half life for this reaction when the starting concentration is 1.12×10-2 M is ___s and when the starting concentration is 5.60×10-3 M is ___s.
(2)The average rate of disappearance of NH3 from t = 0 s to t = 1000 s is ___ M s-1.
(3)The average rate of disappearance of NH3 from t = 1000 s to t = 1.51×103 s is ___ M s-1.
(4)Based on these data, the rate constant for this _______zero,first or second order reaction is ___M s-1.
Answer #1

1) If the initial concentration is 1.12x10 ^ -2 M, the average life time is 1000 s.

If the initial concentration is 5.6x10 ^ -3 M, the average life time is 510 s.

2) The rate is calculated:

Rate = ΔM / Δt = (5.6x10 ^ -3 - 1.12x10 ^ -2) / (1000 - 0) = 5.6x10 ^ -6 M / s

3) The rate is calculated:

Rate = (2.8x10 ^ -3 - 5.6x10 ^ -3) / (1.51x10 ^ 3 - 1000) = 5.5x10 ^ -6 M / s

4) The reaction is zero order, since the reaction rate does not change with respect to the concentration.

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.